OpenStudy (anonymous):

Which fraction in simplest form is equivalent to 0.4? (see reply bar for multiple choice) (im checking my work. I got 4/16 but I don't think that's right :I)

OpenStudy (akashdeepdeb):

0.4 = 4/10 Simplifying 4/10 we get 2/5 Understood? :)

OpenStudy (anonymous):

4/16 = 1/4 = 0.25 so its not 4/16 0.4 = 4/10 = 2/5

OpenStudy (anonymous):

That cant be right because my options are : A. \[\frac{ 4 }{ 16 }\] B. \[\frac{ 80 }{ 32 }\] C.\[\frac{ 12 }{ 3 }\] D. \[\frac{ 24 }{ 60 }\]

OpenStudy (anonymous):

@zarax

OpenStudy (anonymous):

look at your answer choices....which one reduces to 2/5 ?

OpenStudy (anonymous):

A = 1/4 = 0.25 (divide by 4) B = 5/2 = 2.5 (divide by 16) C = 4 (divide by 3) D = 2/5 = 0.4 (divide by 12)

OpenStudy (anonymous):

Oh srry (feels really dumb) thanks